Write four solutions for each of the following equations:

  πx + y = 9

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

To find the four solutions of πx + y = 9 we substitute different values for x and y

x = 0



πx + y = 9

(π × 0)+y = 9

y = 9

(0,9)

Let x = 1

Then,

πx + y = 9

(π×1)+y = 9

π+y = 9

y = 9-π

(1,9-π)

Let y = 0

Then,

πx + y = 9

πx +0 = 9

πx = 9

x =9/π

(9/π,0)

Let x = -1

Then,

πx + y = 9

(π(-1))+y = 9

-π + y = 9

y = 9+π

(-1,9+π)

The solutions are (0,9), (1,9-π),(9/π,0),(-1,9+π)

Thanks